# Antiferromagnetic pseudogap in the two-dimensional Hubbard model deep in the renormalized classical regime

### v1

@@ I Introduction
The pseudogap in the high temperature superconductors was observed in both hole-doped [ding_spectroscopic_1996, loeser_excitation_1996, renner_pseudogap_1998, timusk1999pseudogap, vishik2018photoemission] and electron-doped materials [armitage2001anomalous, motoyama_spin_2007]. Many theories were proposed for explanation of the pseudogap [kampf1990pseudogaps, Vilk1996, Vilk1997, varma1997non, norman1998phenomenology, franz1998phase, chen1998pairing, schmalian_microscopic_1999, senechal2004hot, kyung2006pseudogap, Ferrero_Cornaglia_De_Leo_Parcollet_Kotliar_Georges_2009, Sordi_Haule_Tremblay_2010, Gull_Ferrero_Parcollet_Georges_Millis_2010, sordi2012pseudogap, Gull_Millis_2013, Sordi_Semon_Haule_tremblay_2013, atkinson2015charge, Fratino_Sémon_Sordi_Tremblay_2016, Wu_Scheurer_Chatterjee_Sachdev_Georges_Ferrero_2018, wu2021interplay, Fratino_Bag_Camjayi_Civelli_Rozenberg_2021, gauvin2022disorder, Walsh_Charlebois_Sémon_Sordi_Tremblay_2022, wang2023phase, ye_location_2023, dai2020modeling, Ye_Chubukov_2023, sakai2023nonperturbative]. There is no current consensus on the nature of the pseudogap. One of the discussed mechanisms for the pseudogap is critical antiferromagnetic fluctuations. In the hole-doped materials the antiferromagnetic correlation length is short, it does not grow fast with decreasing temperature and the pseudogap occurs above the superconducting state. It thus seems unlikely that the origin of the pseudogap in the hole-doped high temperature superconductors is due to long wavelength antiferromagnetic fluctuations.
By contrast, in the the electron-doped high temperature superconductors, the pseudogap was observed in the regime where the antiferromagnetic correlation length grows exponentially [armitage2001anomalous, motoyama_spin_2007]. It was shown [Kyung2004, motoyama_spin_2007] that the pseudogap in these quasi-two-dimensional materials can be explained quantitatively by critical antiferromagnetic fluctuations in the renormalized classical regime.
@@
Theoretically, this pseudogap in weak to intermediate coupling has been predicted in the context of the repulsive 2D Hubbard model at half-filling using the two-particle self-consistent approximation (TPSC) [Vilk1996, Vilk1997]. The TPSC approach satisfies the Pauli principle, the local spin and charge sum rules and the Mermin-Wagner theorem and predicts a wide renormalized classical (RC) regime in which the pseudogap exists. That work also established the criterion for the pseudogap: the pseudogap appears when the correlation length $\xi$ for spin fluctuations exceeds the thermal de Broglie wavelength of electrons $\xi_{th}=v_{F}/\pi T$ (in units where $\hbar=1$ and $k_{B}=1$). Here $v_{F}$ is an electron velocity at the Fermi surface and $T$ is the temperature  [^1] This criterion can be modified in the presence of disorder [gauvin2022disorder]. The TPSC approach has been extended to more realistic multi-orbital models [Miyahara_2013, Zantout_2021, Gauvin-Ndiaye_Leblanc_Marin_Martin_Lessnich_Tremblay_2024], to spin-orbit coupling [lessnich2024spin], to out-of-equilibrium calculations [Simard_2022], to the attractive Hubbard model [vilk1998attractive] and to disorder effects [gauvin2022disorder]. The flexibility of this method and its low computational cost makes it attractive for applications to real materials.
@@
Theoretical studies of the antiferromagnetic pseudogap were extended in a number of directions: $t-t^{\prime}$ model in the presence of doping[vilk1997shadow, schmalian_weak_1998, schmalian_microscopic_1999, kuchinskii_models_1999], quasi-two-dimensional system [dare1996crossover, moca_analytical_2000] and approximate methods that go beyond one-loop approximation for the self-energy [sedrakyan_pseudogap_2010, Ye_Chubukov_2023, nikolaenko_spin_2023].
@@
The existence of the pseudogap in the 2D Hubbard model at half-filling has been confirmed by statistically exact numerical calculations with both diagrammatic (DiagMC) [vsimkovic2020extended, Schaefer2021] and determinantal (DQMC) quantum Monte Carlo [Moukouri2000, schafer2015fate, Schaefer2021]. These methods convincingly show that, in this model, the pseudogap develops in the renormalized classical regime. While these numerical methods work well in the metallic state and when the pseudogap begins to open-up, they run into convergence problems at low temperatures when the spin correlation length becomes large. They are also computationally expensive and difficult to generalize to the more realistic three-band models. Hence we go back to the TPSC approach to apply it to regimes that are presently inaccessible to exact methods.
@@
However, while the TPSC approach agrees quantitatively with benchmarks Monte Carlo results at high temperatures [Vilk1997, TremblayMancini:2011, Moukouri2000] and gives a description of the crossover to the renormalized classical regime, it fails deep inside the pseudogap regime where it leads to some nonphysical results. For example, at half-filling in the nearest-neighbor hopping Hubbard model, it predicts that double occupancy decreases to zero at zero temperature. The modified TPSC+ approach proposed in Ref. [gauvin2023improved] remedies this by taking into account the feedback of the single-particle properties on the spin and charge susceptibilities. TPSC+ provides a framework for studying strongly correlated electron systems deep inside the pseudogap regime where no well-defined quasiparticles exist. The approach also improved agreement with benchmarks in the metallic regime. However, due to technical convergence issues, the study in Ref. [gauvin2023improved] was limited to temperatures above the pseudogap crossover defined by DiagMC results.
@@ II.2 The two-particle self-consistent (TPSC) approach
In this subsection, we describe some of the main properties of the TPSC approach. This non-perturbative method satisfies the Pauli principle, the Mermin-Wagner theorem and conservation laws for the spin and charge susceptibilities. The spin and charge susceptibilities have RPA-like forms but with bare interaction $U$ replaced by renormalized spin $U_{sp}$ and charge $U_{ch}$ interactions
@@
The susceptibility $\chi^{(1)}(q)$ is identical to the non-interacting case but it is important to keep in mind that the self-energy entering the Green’s function $\mathcal{G}^{(1)}_{\sigma}(k)$ at this level is a constant which can be absorbed into the chemical potential for a given filling $n$ [TremblayMancini:2011].
The renormalized vertices $U_{sp}$ and $U_{ch}$ are found self-consistently using the local spin and charge sum rules [vilk1994theory]:
@@
To compute self-consistently the vertices $U_{sp}$ and $U_{ch}$ from the sum rules, a third equation is needed. In TPSC, the following ansatz [vilk1994theory, hedayati1989ground] for $U_{sp}$ is used:
@@
The renormalized vertex $U_{sp}$ is reduced relative to the bare interaction by the pair correlation function $g_{\uparrow\downarrow}(0)=\langle n_{\uparrow}n_{\downarrow}\rangle/\langle n_{\uparrow}\rangle\langle n_{\downarrow}\rangle$. The Hartree-Fock approximation corresponds to the non-interacting value $g_{\uparrow\downarrow}(0)=1$ and the reduction of $g_{\uparrow\downarrow}(0)$ from unity describes Kanamori-Brückner screening. The latter is responsible for the absence of the ferromagnetic order in the Hubbard model [vilk1994theory].
In analogy with the procedure in the electron gas, the self-energy is computed in a second step, using the spin and charge susceptibilities from the first step that contain collective modes [Vilk1996, Vilk1997, Moukouri2000]:
@@
In the original TPSC paper [Vilk1996], [Vilk1997] the value $\tilde{g}=1/4$ was used and in a later work [Moukouri2000] it was argued that $\tilde{g}=3/8$ was a better choice. The ambiguity is due to the fact that a constant vertex $U_{sp}$ violates rotational invariance so that one obtains different formulas in the longitudinal and transverse spin channel. The factor $\tilde{g}=3/8$ comes from averaging longitudinal and transverse channels to restore crossing symmetry (rotational invariance). Here we explore both values, $\tilde{g}=1/4$ and $\tilde{g}=3/8$, to find out the best agreement with Monte Carlo benchmarks.
@@
can be used as an accuracy check  [Vilk1997]. In TPSC, this relation is satisfied for the double-occupancy found at the previous level if $\Sigma_{\sigma}^{(2)}$ from Eq. (8) is used with the Green function $\mathcal{G}_{\sigma}$ replaced by $\mathcal{G}^{(1)}_{\sigma}$. The difference (generally a few percent) with the calculation where $\mathcal{G}_{\sigma}^{(2)}$ is used instead, serves as an accuracy check [Vilk1997]. Note that when $\tilde{g}=3/8$, each of the contributions from the three components of the spin fluctuation channels and from the single charge fluctuation channel contribute equally to the Migdal-Galitskii relation. Decreasing $\tilde{g}$ gives more importance to the charge contribution to this relation.
@@ II.3 Pseudogap and renormalized classical regime in TPSC
The self-consistency condition for the spin susceptibility Eq. (5) leads to the suppression of the antiferromagnetic mean field phase transition. This is because a finite $T$ phase transition ($\xi=\infty$) and the large phase space occupied by long-wavelength fluctuations in two dimensions would cause a divergence of the integral on the left-hand side of Eq. (5). Consequently, the phase transition is replaced by a crossover to the renormalized classical regime $\omega_{sp}<T$ that exists in a wide temperature range. Here the characteristic spin-fluctuation frequency $\omega_{sp}\sim\xi^{-2}$ is small because of critical slowing down. In this regime, the correlation length grows exponentially and quickly overcomes the characteristic thermal wave length of electrons $\xi_{th}$. The latter increases only as $v_{F}/T$ for regular points on the Fermi surface [Vilk1996] and as $1/T^{1/2}$ for the Van Hove point $v_{F}=0$ [vilk2023criteria]. When the correlation length becomes $\xi>>\xi_{th}$ a pseudogap opens up in the single-particle spectra. This pseudogap is a precursor to the antiferromagnetic gap in the ground state. The regimes described above where studied in detail and benchmarked in Ref. [Schaefer2021].
Close to the quantum-critical point where antiferromagnetism disappears, the correlation length grows only algebraically with decreasing temperature [Bergeron:2012] so that it does not overcome quickly the thermal de Broglie wavelength. In that regime, a different criterion for the pseudogap based on second derivative of the spectral weight at $\omega=0$ may be preferable [Ye_Chubukov_2023].
While TPSC correctly describes the beginning of the renormalized classical regime and opening of the pseudogap, it fails deep in the pseudogap regime. In that regime, $U_{sp}$ stays just slightly below $U_{crit}=2/\chi^{(1)}(\mathbf{Q},0)$ where $\mathbf{Q}=(\pi,\pi)$. Since $\chi^{(1)}(\mathbf{Q},0)$ diverges as $T\rightarrow 0$ at half-filling in the near-neighbor hopping model, $U_{crit}$ and thus $U_{sp}$ and double occupancy (Eq. (7)) go to zero as well. This is, obviously, nonphysical. To remedy this behavior, it was suggested in Ref. [gauvin2023improved] to replace $\chi^{(1)}(\mathbf{Q},0)$ with the quantity that remains finite as $T\rightarrow 0$. We discuss this approach in the next section.
@@ II.4 TPSC+ approach
All other self-consistency conditions are the same as in the TPSC and $\mathcal{G}_{\sigma}^{(2)}$ is calculated iteratively using the self-energy Eq. (8). The approach is analogous to the pairing approximation (GG0 theory) for the pair susceptibility introduced by Kadanoff and Martin [Kadanoff_Martin_1961, chen2005bcs, Boyack_2018]. This approach partly takes into account self-energy and vertex corrections [Schaefer2021, gauvin2023improved].
It was shown in Ref. [gauvin2023improved] that the above form of $\chi^{(2)}(\mathbf{q},iq_{n})$ remains finite as $T\rightarrow 0$ due to the opening of the pseudogap and that the value of the gap $\Delta$ is given by a modified Stoner criterion with bare $U$ replaced by the renormalized interaction $U_{sp}$, the irreducible vertex in the spin channel.
@@ II.5 TPSC+Classical
Using the classical contribution to the self-energy Eq. (13) to feedback in $\chi^{(2)}$ is sufficient to obtain qualitatively the same results as those obtained in TPSC+. Moreover, we find that the resulting spin susceptibility gives quantitative results very similar to those obtained using the full self-energy Eq. (8) (see next section). This suggests some cancellation between vertex and quantum self-energy corrections. We call this version of the TPSC+ approach, TPSC+Classical.

### v2

@@ I Introduction
The pseudogap in the high temperature superconductors was observed in both hole-doped [1, 2, 3, 4, 5] and electron-doped materials [6, 7]. Many theories were proposed for explanation of the pseudogap [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. There is no current consensus on the nature of the pseudogap. One of the discussed mechanisms for the pseudogap is critical antiferromagnetic fluctuations. In the hole-doped materials the antiferromagnetic correlation length is short, it does not grow fast with decreasing temperature and the pseudogap occurs above the superconducting state. It thus seems unlikely that the origin of the pseudogap in the hole-doped high temperature superconductors is due to long wavelength antiferromagnetic fluctuations.
By contrast, in the the electron-doped high temperature superconductors, the pseudogap was observed in the regime where the antiferromagnetic correlation length grows exponentially [6, 7]. It was shown [36, 7] that the pseudogap in these quasi-two-dimensional materials can be explained quantitatively by critical antiferromagnetic fluctuations in the renormalized classical regime.
@@
Theoretically, this pseudogap in weak to intermediate coupling has been predicted in the context of the repulsive 2D Hubbard model at half-filling using the two-particle self-consistent approximation (TPSC) [9, 10]. The TPSC approach satisfies the Pauli principle, the local spin and charge sum rules and the Mermin-Wagner theorem and predicts a wide renormalized classical (RC) regime in which the pseudogap exists. That work also established the criterion for the pseudogap: the pseudogap appears when the correlation length $\xi$ for spin fluctuations exceeds the thermal de Broglie wavelength of electrons $\xi_{th}=v_{F}/\pi T$ (in units where $\hbar=1$ and $k_{B}=1$). Here $v_{F}$ is an electron velocity at the Fermi surface and $T$ is the temperature  [^1] This criterion can be modified in the presence of disorder [29]. The TPSC approach has been extended to more realistic multi-orbital models [38, 39, 40], to spin-orbit coupling [41], to out-of-equilibrium calculations [42], to the attractive Hubbard model [43] and to disorder effects [29]. The flexibility of this method and its low computational cost makes it attractive for applications to real materials.
@@
Theoretical studies of the antiferromagnetic pseudogap were extended in a number of directions: $t-t^{\prime}$ model in the presence of doping[44, 45, 15, 46], quasi-two-dimensional system [47, 48] and approximate methods that go beyond one-loop approximation for the self-energy [49, 34, 50].
@@
The existence of the pseudogap in the 2D Hubbard model at half-filling has been confirmed by statistically exact numerical calculations with both diagrammatic (DiagMC) [51, 52] and determinantal (DQMC) quantum Monte Carlo [53, 54, 52]. These methods convincingly show that, in this model, the pseudogap develops in the renormalized classical regime. While these numerical methods work well in the metallic state and when the pseudogap begins to open-up, they run into convergence problems at low temperatures when the spin correlation length becomes large. They are also computationally expensive and difficult to generalize to the more realistic three-band models. Hence we go back to the TPSC approach to apply it to regimes that are presently inaccessible to exact methods.
@@
However, while the TPSC approach agrees quantitatively with benchmarks Monte Carlo results at high temperatures [10, 55, 53] and gives a description of the crossover to the renormalized classical regime, it fails deep inside the pseudogap regime where it leads to some nonphysical results. For example, at half-filling in the nearest-neighbor hopping Hubbard model, it predicts that double occupancy decreases to zero at zero temperature. The modified TPSC+ approach proposed in Ref. [56] remedies this by taking into account the feedback of the single-particle properties on the spin and charge susceptibilities. TPSC+ provides a framework for studying strongly correlated electron systems deep inside the pseudogap regime where no well-defined quasiparticles exist. The approach also improved agreement with benchmarks in the metallic regime. However, due to technical convergence issues, the study in Ref. [56] was limited to temperatures above the pseudogap crossover defined by DiagMC results.
@@ II.2 The two-particle self-consistent (TPSC) approach
In this subsection, we describe some of the main properties of the TPSC approach. This nonperturbative method satisfies the Pauli principle, the Mermin-Wagner theorem and conservation laws for the spin and charge susceptibilities. The spin and charge susceptibilities have RPA-like forms but with bare interaction $U$ replaced by renormalized spin $U_{sp}$ and charge $U_{ch}$ interactions
@@
The susceptibility $\chi^{(1)}(q)$ is identical to the noninteracting case but it is important to keep in mind that the self-energy entering the Green’s function $\mathcal{G}^{(1)}_{\sigma}(k)$ at this level is a constant which can be absorbed into the chemical potential for a given filling $n$ [55].
The renormalized vertices $U_{sp}$ and $U_{ch}$ are found self-consistently using the local spin and charge sum rules [57]:
@@
To compute self-consistently the vertices $U_{sp}$ and $U_{ch}$ from the sum rules, a third equation is needed. In TPSC, the following ansatz [57, 58] for $U_{sp}$ is used:
@@
The renormalized vertex $U_{sp}$ is reduced relative to the bare interaction by the pair correlation function $g_{\uparrow\downarrow}(0)=\langle n_{\uparrow}n_{\downarrow}\rangle/\langle n_{\uparrow}\rangle\langle n_{\downarrow}\rangle$. The Hartree-Fock approximation corresponds to the noninteracting value $g_{\uparrow\downarrow}(0)=1$ and the reduction of $g_{\uparrow\downarrow}(0)$ from unity describes Kanamori-Brückner screening. The latter is responsible for the absence of the ferromagnetic order in the Hubbard model [57].
In analogy with the procedure in the electron gas, the self-energy is computed in a second step, using the spin and charge susceptibilities from the first step that contain collective modes [9, 10, 53]:
@@
In the original TPSC paper [9], [10] the value $\tilde{g}=1/4$ was used and in a later work [53] it was argued that $\tilde{g}=3/8$ was a better choice. The ambiguity is due to the fact that a constant vertex $U_{sp}$ violates rotational invariance so that one obtains different formulas in the longitudinal and transverse spin channel. The factor $\tilde{g}=3/8$ comes from averaging longitudinal and transverse channels to restore crossing symmetry (rotational invariance). Here we explore both values, $\tilde{g}=1/4$ and $\tilde{g}=3/8$, to find out the best agreement with Monte Carlo benchmarks.
@@
can be used as an accuracy check  [10]. In TPSC, this relation is satisfied for the double-occupancy found at the previous level if $\Sigma_{\sigma}^{(2)}$ from Eq. (8) is used with the Green function $\mathcal{G}_{\sigma}$ replaced by $\mathcal{G}^{(1)}_{\sigma}$. The difference (generally a few percent) with the calculation where $\mathcal{G}_{\sigma}^{(2)}$ is used instead, serves as an accuracy check [10]. Note that when $\tilde{g}=3/8$, each of the contributions from the three components of the spin fluctuation channels and from the single charge fluctuation channel contribute equally to the Migdal-Galitskii relation. Decreasing $\tilde{g}$ gives more importance to the charge contribution to this relation.
@@ II.3 Pseudogap and renormalized classical regime in TPSC
The self-consistency condition for the spin susceptibility Eq. (5) leads to the suppression of the antiferromagnetic mean field phase transition. This is because a finite $T$ phase transition ($\xi=\infty$) and the large phase space occupied by long-wavelength fluctuations in two dimensions would cause a divergence of the integral on the left-hand side of Eq. (5). Consequently, the phase transition is replaced by a crossover to the renormalized classical regime $\omega_{sp}<T$ that exists in a wide temperature range. Here the characteristic spin-fluctuation frequency $\omega_{sp}\sim\xi^{-2}$ is small because of critical slowing down. In this regime, the correlation length grows exponentially and quickly overcomes the characteristic thermal wavelength of electrons $\xi_{th}$. The latter increases only as $v_{F}/T$ for regular points on the Fermi surface [9] and as $1/T^{1/2}$ for the Van Hove point $v_{F}=0$ [59]. When the correlation length becomes $\xi>>\xi_{th}$ a pseudogap opens up in the single-particle spectra. This pseudogap is a precursor to the antiferromagnetic gap in the ground state. The regimes described above where studied in detail and benchmarked in Ref. [52].
Close to the quantum-critical point where antiferromagnetism disappears, the correlation length grows only algebraically with decreasing temperature [60] so that it does not overcome quickly the thermal de Broglie wavelength. In that regime, a different criterion for the pseudogap based on second derivative of the spectral weight at $\omega=0$ may be preferable [34].
While TPSC correctly describes the beginning of the renormalized classical regime and opening of the pseudogap, it fails deep in the pseudogap regime. In that regime, $U_{sp}$ stays just slightly below $U_{crit}=2/\chi^{(1)}(\mathbf{Q},0)$ where $\mathbf{Q}=(\pi,\pi)$. Since $\chi^{(1)}(\mathbf{Q},0)$ diverges as $T\rightarrow 0$ at half-filling in the near-neighbor hopping model, $U_{crit}$ and thus $U_{sp}$ and double occupancy (Eq. (7)) go to zero as well. This is, obviously, nonphysical. To remedy this behavior, it was suggested in Ref. [56] to replace $\chi^{(1)}(\mathbf{Q},0)$ with the quantity that remains finite as $T\rightarrow 0$. We discuss this approach in the next section.
@@ II.4 TPSC+ approach
All other self-consistency conditions are the same as in the TPSC and $\mathcal{G}_{\sigma}^{(2)}$ is calculated iteratively using the self-energy Eq. (8). The approach is analogous to the pairing approximation (GG0 theory) for the pair susceptibility introduced by Kadanoff and Martin [61, 62, 63]. This approach partly takes into account self-energy and vertex corrections [52, 56].
It was shown in Ref. [56] that the above form of $\chi^{(2)}(\mathbf{q},iq_{n})$ remains finite as $T\rightarrow 0$ due to the opening of the pseudogap and that the value of the gap $\Delta$ is given by a modified Stoner criterion with bare $U$ replaced by the renormalized interaction $U_{sp}$, the irreducible vertex in the spin channel.
@@ II.5 TPSC+Classical
Using the classical contribution to the self-energy Eq. (13) to feedback in $\chi^{(2)}$ is sufficient to obtain qualitatively the same results as those obtained in TPSC+. Moreover, we find that the resulting spin susceptibility gives quantitative results very similar to those obtained using the full self-energy Eq. (8) (see next section). This suggests some cancellation between vertex and quantum self-energy corrections. We call this version of the TPSC+ approach, TPSC+Classical. It uses minimal necessary feedback self-energy to get good description of the spin susceptibility deep in the renormalized classical regime. The use of momentum and frequency dependent feedback self-energy together with constant vertex involves trade off. It gives better description of the antiferromagnetic fluctuation but it is not consistent with conservation laws. It thus make sense to keep the feedback self-energy in the calculation of the susceptibility to a minimum.
@@ III.1 Comparison with DiagMC results
III.1.1 Spin Susceptibility
We start with results for the maximum of spin susceptibility as a function of the interaction strength. Fig. 1 shows $\chi_{sp}(\mathbf{q}_{max},0)$ as a function of $U$ for $n=0.875$ and $T=0.2$. In this figure TPSC+Classical-1 uses $\tilde{g}_{fb}=1/4$ for feedback self-energy and TPSC+Classical-2 uses $\tilde{g}_{fb}=3/8$. The TPSC+ version uses full self-energy Eq. (8) with both spin and charge contributions and $\tilde{g}_{fb}=3/8$ while TPSC+SFM uses only spin contribution with $\tilde{g}_{fb}=1/4$. As we can see from figure Fig. 1, TPSC+Classical-1 has almost the same results as TPSC+SFM while TPSC+Classical-2 has almost the same results as TPSC+ for $U\leq 4$. This implies that quantum contributions to the feedback self-energy has rather small effect on the spin susceptibility $\chi_{sp}(\mathbf{q}_{max},0)$ for these values of $U$. We found similar results for other fillings $n$ and temperatures $T$ and $U\leq 4$. This suggests some cancellation between self-energy and vertex corrections. The area of the applicability for all TPSC+ variants is weak to intermediate interaction regime $U\leq 4$. Comparison with Monte Carlo results [65] in the Fig. 1 shows that for $U>4$ the Monte Carlo values of $\chi_{sp}(\mathbf{q}_{max},0)$ start to saturate and then decline as a function of $U$. The latter behavior indicates transition to the Mott physics (localized spins) in the strong coupling limit. All TPSC+ variants miss this behavior quantitatively. We note, however, that all TPSC+ variants do have a maximum of $\chi_{sp}(\mathbf{q}_{max},0)$ as a function of $U$ but it occurs at larger $U\geq 6$ and it is much broader than Monte Carlo result. Nevertheless, it is interesting because in the original TPSC theory $\chi_{sp}(\mathbf{q}_{max},0)$ saturates with increasing $U$ but does not decrease. Thus the decrease of $\chi_{sp}(\mathbf{q}_{max},0)$ at large $U$ in our theory is solely due to the effect of the feedback self-energy on the spin susceptibility. In the opposite limit of relatively small $U\leq 2.5$, all TPSC+ variants agree quantitatively with Monte Carlo data.